OpenStudy (erniellyd28):

a/b-c/b=1 (solve the formula for b)

OpenStudy (anonymous):

\[\frac{ a }{ b }-\frac{ c }{ b }=1?\]

OpenStudy (erniellyd28):

yes

OpenStudy (anonymous):

than \[\frac{ a-c }{ b }=1\] \[b=a-c\]

OpenStudy (erniellyd28):

thank you very much would it be ok if we did 2 more please?

OpenStudy (anonymous):

I can try

OpenStudy (erniellyd28):

r=2k+2u(solve for k)

OpenStudy (anonymous):

r = 2k + 2u r -2u = 2k \[\frac{ r-2u }{ 2 }=k\]

OpenStudy (erniellyd28):

im learning something new i dont know why i can never get math

OpenStudy (ihelpyoulearn):

Still need help with this?

OpenStudy (erniellyd28):

you mean on this problem? or another one?

OpenStudy (ihelpyoulearn):

Any :)

OpenStudy (erniellyd28):

one more then i will let you go :-)

OpenStudy (ihelpyoulearn):

I can try once I see it

OpenStudy (ihelpyoulearn):

Or is it already posted?

OpenStudy (erniellyd28):

no not yet

OpenStudy (erniellyd28):

c=1/2 pie n squared x

OpenStudy (anonymous):

\[c=\frac{ 1 }{ 2 }\times \pi \times n ^{x} ?\]

OpenStudy (anonymous):

be back

OpenStudy (erniellyd28):

take off the first x and move the n squared to the side of the second x

OpenStudy (anonymous):

x is time not letter x

OpenStudy (erniellyd28):

i was trying to draw it for you but it would not let me

OpenStudy (anonymous):

\[c=\frac{ 1 }{2 }*\pi*n ^{x} ?\]

OpenStudy (erniellyd28):



OpenStudy (anonymous):